OpenStudy (tiffany_rhodes):

Determine the trajectories for the nonlinear system, dx/dt = y and dy/dt = x + 13x^3, by integrating the equation for dy/dx

OpenStudy (irishboy123):

dy/dx = (dy/dt) / (dx/dt) dy/dx = (x + 13 x^3 ) / y y dy = (x + 13 x^3 ) dx ∫ y dy = ∫ (x + 13 x^3 ) dx ok?

OpenStudy (tiffany_rhodes):

yeah, thanks!